# Explicit core-hole single-particle methods for - and -edge X-ray absorption and electron energy-loss spectra

### v1

@@ IV.2 Solids
We then consider solids, which we describe in a bulk approximation without the consideration of a surface. We first examine the Ti edge or the Ti 2 core-transitions. We analyze three different solid systems that feature a Ti 2 core hole in the following: SrTiO3, the rutile structure of TiO2, and the anatase structure of TiO2.
The unit cell of SrTiO3 has a structure. The initial structure was obtained from the Materials Project (MP)[77] with ID: mp-5229 and subsequently re-relaxed in a supercell. The Brillouin zone was sampled using three -points in each direction. The Ti 2 XAS spectrum of SrTiO3, has been investigated experimentally[75] and theoretically[34, 78, 79, 80] in great detail in the literature. The spectra are presented in Fig. 8, where the experimental spectrum[75] consists of four main peaks with varying widths. Both TP and XCH qualitatively reproduce this general peak structure.
@@
We next turn to TiO2 with its most common rutile and anatase crystal structures. The TiO2 rutile structure belongs to the space group, where the initial structure from MP with ID mp-2657 was re-relaxed in the supercell. The Brillouin zone was sampled by five -points in each direction. The TiO2 anatase structure is part of the space group, where the initial structure from MP with ID mp-390 was re-relaxed in the supercell. The Brillouin zone was also sampled by five -points in each direction. The calculated spectra for rutile are shown in Fig. 9 a) and b), while c) displays an experimental spectrum reported in Ref. 76. In Fig. 9 d), e), and f), we present similar data for the anatase structure, using the experimental spectra from the same study.
@@
We now consider NiO, an antiferromagnetic solid that exhibits a rocksalt crystal structure with a lattice constant of [77, 83]. We use a cubic supercell and sample the Brillouin zone with five -points in each direction.
@@
When comparing the two theoretical spectra to the experimental spectra in d), we notice that they are both in quite good agreement. They reproduce the general shape of the spectra; however, the proportions in the XCH in b) seem exaggerated, while the TP method offers a more accurate representation. This is also reflected in the R2 values. The comparison between the theoretical spectra and the experimental results in c) can be found in Sec. VI in SI.
@@
We also wish to address the NEXAFS spectra, specifically concerning -type core holes. Here, we have selected tin as element and compared our calculations to the available experimental pectrum of stannic oxide (SnO2). There are numerous gas-phase XPS spectra for Sn 3, making the fit for the corresponding absolute shift unproblematic, as shown in Fig. S3 in SI. Unfortunately, we have not found any molecular NEXAFS spectrum for comparison. Thus, we are limited to the bulk phase, excluding TDDFT.
SnO2 exhibits a rutile crystal structure[85] (space group 42/) with lattice constants 4.76 Å, 3.21 Å. [77]. The initial structure was sourced from the Materials Projects SnO2, with the materials ID: mp-856.[77]. Our calculation is done with a supercell with six -points in each direction; however, the calculations converge with four -points. Further details can be found in Secs. VII and VI in SI.
Fig. 11 shows the spectra produced by the TP method with in a), TP with in b), and the XCH method in c). The experimental spectra by Chuvenkova et al. [84] is shown in d). All theoretical spectra consistently reproduce the features of the experiment in d) with quite high R2 values, with the TP () and XCH methods demonstrating the best agreement. However, the XCH spectrum shows the smallest energy shift compared to the experiment.
@@ IV.3 Electron energy-loss in graphene
Alignment of the first prominent peaks around 105 eV leads to a small mismatch of the broad peaks around 130 eV, where in particular the XCH is at slightly lower energy. This suggests and insufficient screening of the full core-hole in XCH, compared to TP with . Indeed, the reduction of the core-hole to helps in this respect, leading to very good agreement for both binding situations. The predicted spectra indeed allow to distinguish the binding mode, as is explicitly shown in Fig. S7 in SI. Overall, it is remarkable how well the simulations are able to reproduce the EELS fine structure, as has also been observed previously without consideration of an explicit core-hole[29, 30, 31, 11].
@@ V Conclusions
Both the XCH and TP methods accurately replicate the -edge XANES experimental spectra for the majority of the molecular systems examined. Their performance is comparable to the LrTDDFT method and, in some instances, even surpasses it. In our comparison of the TP and XCH methods for molecular systems, we observed similar R2 values, with the XCH method being slightly higher in Si cases and the TP method in S cases. For solid systems, TP slightly outperformed XCH in most instances, except for SnO2, where XCH achieved better results. The comparison with experimental EEL spectra is improved when the occupation of the core-hole is further reduced than in the traditional TP using a half occupied core-hole, however.
@@
Overall, the explicit core-hole TP and XCH methods are able to sufficiently describe -edge XANES for the majority of systems we evaluated. There is no clear preference for either of the two methods as their advantages and disadvantages seem to be system-dependent. Nevertheless, the methods described represent a computationally cheap method to simulate XAS spectra with an overall good agreement to experiment on the absolute energy scale.
@@ Acknowledgements.
E. J. and M. W. thank L. Mayrhofer and M. Moseler for useful discussion and for reading the manuscript. E. J., N. H. and M. W. acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Excellence Cluster LivMatS under Germany’s Excellence Strategy—EXC-2193/1— 390951807. E.J. and M.W. are thankful for the computing resources provided by the state of Baden-Württemberg through bwHPC and the German Research Foundation through grant number INST 40/575-1 FUGG (NEMO and JUSTUS2 clusters). We thank Quentin Ramasse for providing the raw data for the experimental EEL spectra considered in this study.VI Author declarations
VI.1 Conflict of Interest
@@ VI.2 Author Contributions
Data is available on request from the authors. The code developed to perform these calculations will soon be available in the GPAW library.

### v2

@@ IV.2 Solids
We then consider solids, which we describe in a bulk approximation without the consideration of a surface. We solely use single-electron methods for solids due to the limitations of LrDDFT in the application to solids.[26, 93] We first examine the Ti edge or the Ti 2 core-transitions. We analyze three different solid systems that feature a Ti 2 core hole in the following: SrTiO3, the rutile structure of TiO2, and the anatase structure of TiO2.
The unit cell of SrTiO3 has a structure. The initial structure was obtained from the Materials Project (MP)[96] with ID: mp-5229 and subsequently re-relaxed in a supercell. The Brillouin zone was sampled using three -points in each direction within the supercell. The Ti 2 XAS spectrum of SrTiO3, has been investigated experimentally[94] and theoretically[37, 97, 98, 99] in great detail in the literature. The spectra are presented in Fig. 9, where the experimental spectrum[94] consists of four main peaks with varying widths. Both TP and XCH qualitatively reproduce this general peak structure.
@@
We next turn to TiO2 with its most common rutile and anatase crystal structures. The TiO2 rutile structure belongs to the space group, where the initial structure from MP with ID mp-2657 was re-relaxed in the supercell. The Brillouin zone was sampled by five -points in each direction within the supercell. The TiO2 anatase structure is part of the space group, where the initial structure from MP with ID mp-390 was re-relaxed in the supercell. The Brillouin zone was also sampled by five -points in each direction within the supercell. The calculated spectra for rutile are shown in Fig. 10 a) and b), while c) displays an experimental spectrum reported in Ref. 95. In Fig. 10 d), e), and f), we present similar data for the anatase structure, using the experimental spectra from the same study.
@@
We now consider NiO, an antiferromagnetic solid that exhibits a rocksalt crystal structure with a lattice constant of [96, 102]. We use a cubic supercell and sample the Brillouin zone with five -points in each direction within the supercell.
@@
When comparing the two theoretical spectra to the experimental spectra in d), we notice that they are both in quite good agreement. They reproduce the general shape of the spectra; however, the proportions in the XCH in b) seem exaggerated, while the TP method offers a more accurate representation. This is also reflected in the R2 values. The comparison between the theoretical spectra and the experimental results in c) can be found in Sec. VI in SM.
@@
We also wish to address the NEXAFS spectra, specifically concerning -type core holes. Here, we have selected tin as element and compared our calculations to the available experimental spectrum of stannic oxide (SnO2). There are numerous gas-phase XPS spectra for Sn 3, making the fit for the corresponding absolute shift unproblematic, as shown in Fig. S3 in SM. Unfortunately, we have not found any molecular NEXAFS spectrum for comparison. Thus, we are limited to the bulk phase, excluding TDDFT.
SnO2 exhibits a rutile crystal structure[104] (space group 42/) with lattice constants 4.76 Å, 3.21 Å. [96]. The initial structure was sourced from the Materials Projects SnO2, with the materials ID: mp-856.[96]. Our calculation is done with a supercell with six -points in each direction within the supercell; however, the calculations converge with four -points. Further details can be found in Secs. VII and VI in SM.
Fig. 12 shows the spectra produced by the TP and XCH compared to the experimental spectra by Chuvenkova et al. [103]. Both theoretical spectra consistently reproduce the features of the experiment in a) with quite high R2 values, with the XCH method demonstrating the best agreement. The XCH spectrum also has the smallest energy shift compared to the experiment.
@@ IV.3 Electron energy-loss in graphene
Alignment of the first prominent peaks around 105 eV leads to a small mismatch of the broad peaks around 130 eV, where in particular the XCH is at slightly lower energy. This suggests and insufficient screening of the full core-hole in XCH compared to TP with . Indeed, the further reduction of the core-hole charge to helps in this respect, leading to very good agreement for both binding situations. The predicted spectra indeed allow to distinguish the binding mode, as is explicitly shown in Fig. S7 in SM. Overall, it is remarkable how well the simulations are able to reproduce the EELS fine structure, as has also been observed previously without consideration of an explicit core-hole[32, 33, 34, 11].
@@ V Conclusions
Both the XCH and TP methods accurately replicate the -edge XANES experimental spectra for the majority of the molecular systems examined. Including the semi-empirical shifts, the spectra provided are on the absolute energy scale with deviations mostly below 1 eV. These asre in the range of possible experimental uncertainties as the comparison of experimental data for the same molecules shows. The performance of TP and XCH is comparable to the LrTDDFT method and, in some instances, even surpasses it. In our comparison of the TP and XCH methods for molecular systems, we observed similar R2 values, with the XCH method being slightly higher in Si cases and the TP method in S cases. For solid systems, TP slightly outperformed XCH in most instances, except for SnO2, where XCH achieved better results. The comparison with experimental EEL spectra is improved when the occupation of the core-hole is further reduced than in the traditional TP using a half occupied core-hole, however.
@@
Overall, the explicit core-hole TP and XCH methods are able to sufficiently describe -edge XANES for the majority of systems we evaluated. There is no clear preference for either of the two methods as their advantages and disadvantages seem to be system-dependent. Nevertheless, the methods described represent a computationally cheap way (a speed up factor of 40 relative to TDDFT) to simulate XAS spectra with an overall good agreement to experiment on the absolute energy scale. This may be very useful for a fast high-throughput screening of spectra from a large group of materials as was applied e.g. for Raman spectra[106].
@@
VI Supplementary Material
A comparison of the XPS and XAS nomenclature. Details on the implementation, the determination of semiempirical shifts and the spin-orbit coupling calculations. List of the constants used. Details about the optimal fitting of the spectra. Convergence tests.
@@ Acknowledgements.
E. J. and M. W. thank L. Mayrhofer and M. Moseler for useful discussion and for reading the manuscript. E. J., N. H. and M. W. acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Excellence Cluster LivMatS under Germany’s Excellence Strategy—EXC-2193/1— 390951807. E.J. and M.W. are thankful for the computing resources provided by the state of Baden-Württemberg through bwHPC and the German Research Foundation through grant number INST 40/575-1 FUGG (NEMO and JUSTUS2 clusters). We thank Quentin Ramasse for providing the raw data for the experimental EEL spectra considered in this study.VII Author declarations
VII.1 Conflict of Interest
@@ VII.2 Author Contributions
Data is available on request from the authors. The code developed to perform these calculations is available in the GPAW library.